# Tunable transport in bi-disperse porous materials with vascular structure

### v2

@@ I Introduction
As we have presented in previous studies, synthetic systems inspired by the structure and hypothesized function of xylem provide a route to generating extreme scenarios in which the pore liquid experiences large tensions and to studying imbibition and permeation dynamics driven by these stresses [13, 14, 15]. Such synthetic xylem structures provide a basis for testing biophysical hypotheses [13, 16], interrogating nanoscale transport and thermodynamic phenomena [17], and designing new approaches to heat transfer [18, 19] and separations [20].
@@ II.1 Samples
Table 1: Sample properties: dimensions of microchannel segments ($w$, $\ell$, $d$) as defined in Figure 2, and corresponding aspect ratio ($\xi$) and area coverage ratio ($\gamma$); $\kappa/\kappa_{0}$ is the permeability increase predicted by our numerical simulations (see section III), or by our approximate analytical formula (Equation 17), except for the Blank sample for which $\kappa/\kappa_{0}=1$ by definition.
@@
We fabricated five different geometries for the microstructure: a reference sample (Blank) without any etched microchannels in the glass, serving as a reference for the measure of transport in the bare porous silicon layer (permeability $\kappa_{0}\,\mathrm{[m^{2}/(Pa.s)]}$), and four structures of similar surface coverage ($\gamma=S_{\mathrm{microchannels}}/S_{\mathrm{poSi}}\mathrm{[-]}$, where $S_{\mathrm{microchannels}}\,\mathrm{[m^{2}]}$ is the in-plane area of all microchannels and $S_{\mathrm{poSi}}$ is the total in-plane area of the porous silicon through which transport occurred), but strongly varying in-plane aspect ratio ($\xi\mathrm{[-]}$). Table 1 provides the geometries of the features of the five samples; see also Supplemental Material, Figure 9 for actual micrographs of every sample. For all samples, the structure was repeated across a zone of length $L=1$ cm and width $W=2.2$ cm, except for the vein sample with $L=1.5$ cm. As a result, each vein in this latter sample spanned almost the whole length of the medium, while for other samples the pattern was repeated in the $y$ direction.
@@ II.2 Experiments
For high pressure flow experiments (Figure 3a), we first removed air from the pores and microchannels by evacuation in vacuum. Then, we placed the samples in a commercial, high-pressure vessel (HIP Inc.) 111 Note that appropriate equipment and protection must be used when handling fluids at high pressures. filled with water, with an applied pressure, $P$ (typically 3 to 20 MPa depending on the sample). This high pressure resulted in an inwards flow, progressively filling the serpentine channel and shrinking the bubble at the backend of the sample (bottom of structures in Figure 2). Samples were evacuated prior to filling to avoid the pressure associated with compressing and dissolving trapped air. We left the samples under pressure for intervals of time, $\Delta t$ before releasing the pressure. We took snapshots of the samples out of the vessel with a stereoscope and camera (Leica MZFLIII stereoscope from Leica Microsystems GmbH, Wetzlar, Germany, and a QImaging Retiga 1300 camera from QImaging, Surrey, Canada), in order to measure the displacement, $\Delta X$ of the meniscus in the serpentine channel (see Supplemental Material, Figure 10); we repeated this process until filling was complete. We estimated the velocity for each period of pressurization, $U=\Delta X/\Delta t$. This process was required because the pressure vessel was opaque. We then estimated the mass flow rate, Q [kg/s] as:
@@
We note that in some experiments, the meniscus was in the distribution channel rather than in the serpentine channel (see e.g. Supplemental Material, Figure 10). In both cases, we used the known cross-sectional area of the channel, $s\,\mathrm{[m^{2}]}$, in which we tracked the meniscus to convert the speed of the meniscus into a volumetric flow rate.
@@
In this scenario, inspired by xylem sap flow in plants, flow through the structure was driven by evaporation at the open edge (the "leaf") into a subsaturated vapor (Figure 3b). In a previous publication [17], we showed that this method provided an accurate way of measuring permeability of nanoporous samples. Here, we apply it to the composite, synthetic xylem structures. In order to perform transpiration experiments, we used samples that were completely filled with water except for a remaining vapor bubble in the serpentine channel. This situation occurred naturally at the end of a high pressure experiment (see above). We then placed the samples in a humidity-controlled environment with sub-saturated water vapor (water vapor pressure, $p<p_{\mathrm{sat}}(T)$). Due to the local equilibrium at the open edge between liquid water in the nanopores and the external water vapor (see Theory section), a reduced (typically negative) pressure developed at the open edge, resulting, after a short transient, in a steady-state flow through the structure [17]. This flow progressively emptied the serpentine channel; we tracked the resulting motion of the liquid/vapor meniscus in the serpentine channel, using a camera (Point Grey Grasshopper Monochrome camera) a macro lens (AF Micro-Nikkor 60 mm f/2.8), and white, diffuse illumination (Schott ACE 150W light source). From the velocity of the meniscus, we calculated the mass flow rate through the structure using Equation (1). For every sample we recorded transpiration-induced flow at a variety of imposed vapor pressures, $p$, in a range high enough to avoid triggering either dewetting (desorption) from the nanopores or cavitation within the microchannels (see Results section below).
@@
We used spontaneous imbibition as a third method to drive transport within our structures. As we have shown recently [21], liquid imbibition in a nanoporous medium can occur spontaneously by condensation of the vapor into the nanopores, provided that the vapor pressure of the external vapor is above the threshold for capillary condensation, $p_{\mathrm{cond}}$. With our poSi layer, $p_{\mathrm{cond}}\simeq 0.6\,p_{\mathrm{sat}}$ [21], so that we worked with an imposed vapor pressure above this value, $p=0.93\,p_{\mathrm{sat}}$ to induce condensation and imbibition. We followed the imbibition dynamics optically, using time-lapse videos obtained with the same imaging system as for transpiration experiments (camera and macro lens, see above). Because the optical reflectance of the poSi layers changes as a function of water content (see Results), we tracked the imbibition front by monitoring relative changes in the gray level of the image, $\Delta I/I$ averaged in the $x$ direction; we then extracted the position of the imbibition front by defining $Y(t)$ as the position at which the signal is at its midpoint between the dry gray level value and the wet value (see [21] for details). From the displacement, $Y(t)$ of the invading liquid front within the structure, we measured the dynamics of imbibition in our structured samples and estimated the effective permeability of the sample using a modified Lucas-Washburn equation (see Theory).
We performed two types of imbibition experiments. For the first type, we first dried the samples for at least 48 hours in dry air (humidity $6$% RH). Then, we placed the samples in a closed box with a humidity of $93$% RH, set by an unsaturated solution of sodium chloride (NaCl). For the second type, the samples were evacuated for at least 24 hours in a vacuum chamber. Then, we released water vapor in the chamber to impose a fixed vapor pressure of $p=0.93\,p_{\mathrm{sat}}$ ($93$% RH). As a result, imbibition occurred with air-filled nanopores and microchannels in the first type of experiments and with evacuated nanopores and microchannels in the second type.
@@ III.3 Relative magnitudes of transport mechanisms
$\frac{q_{\mathrm{v,vac}}}{q_{\mathrm{po}}}=\frac{(\epsilon aH_{\mathrm{v}})^{2}}{12\eta^{\ast}\kappa}=\mathcal{O}(10^{2})$ (10)
@@
$\frac{q_{\mathrm{v,air}}}{q_{\mathrm{po}}}=\frac{\epsilon^{2}aD}{p_{\mathrm{sat}}\kappa}=\mathcal{O}(10^{-1})$ (11)
@@ III.4 Effective permeability of the composite structures
Figure 5: Numerical and analytical predictions of steady-state flow through the composite matrix, as a function of the microchannel aspect ratio ($\xi=\ell/w$) and of the surface density of channels ($\gamma$). (a) Pressure field (colormap) and streamlines (black lines) for the unit cell of the composite porous medium, obtained from numerical simulations (here with $\gamma=0.3$, $\xi=2.2$); inside the white dotted line is the microchannel over the porous layer and outside this line is just the porous substrate. In this prediction, we assumed that the microchannels had infinite permeability, a reasonable approximation when the microchannels are filled with liquid water or pure water vapor (see Theory, Equations (9) and (10)). (b) Comparison of numerical and analytical predictions of the effective permeability ratio $\kappa/\kappa_{0}$ as a function of $\xi$ for different values of $\gamma$. Symbols represent predictions of the numerical simulations as in (a), and lines represent approximate, analytical predictions based on a resistors-in-series approach (Equation 18). Sketches show the appearance of a unit cell for the three different values of $\gamma$, at an aspect ratio, $\xi=8$; the data corresponding to these three sketches is circled in gray (dashes). See Figure 11 in the Supplemental Material for a similar comparison between simulation results and a Maxwell-Garnett-like approach (Equation 16).
@@
(details of the derivation can be found in the Supplemental Material, and in [30]). Equation (16) gave satisfactory agreement with our simulation results, but only for low areal fractions, $\gamma$, and moderate aspect ratios, $\xi$ (see Supplemental Material, Figure 11). This observation is not surprising, given that Zimmerman’s calculation is a first-order approach which performs best at low densities of inclusions [32]. As a result, we also used another approach better suited for large areal coverages and large aspect ratios.
@@ IV.1 High pressure flow
We used high pressure, steady-state flow in our various composite structures to estimate the permeability associated with samples containing microstructure (samples Square, Short, Long, Vein), relative to that of a purely nanoporous sample (Blank, permeability $\kappa_{0}$). From the observed filling velocity $U=dX/dt$ of the meniscus in the serpentine channel (see Methods and Supplemental Material, Figure 10), and Equations (1) and (19) we estimated the permeability ratio $\kappa/\kappa_{0}$, where the index 00 refers to the reference sample (Blank). We report and discuss the extracted values of the permeability ratios in the general discussion (see below, section IV.4 and Figure 8b).
@@ IV.3 Imbibition
Figure 7: Dynamics of spontaneous imbibition by capillary condensation (see Figure 3c). (a) Images recorded during the experiments in the Long sample when evacuated. The darker zone on top of the images is the area wetted by the fluid (contrast has been enhanced for clarity). See Supplemental Material for the corresponding complete video. (b-c) Dynamics of imbibition for all samples when evacuated (b) and air-filled (c), represented by the squared position, $Y^{2}$, of the wetting front as a function of time (see Figure 3c). Different colors represent the different sample geometries, and black lines represent linear fits, indicating Lucas-Washburn-like dynamics on average (Equation 22).
@@
Noticeably, the Vein sample displayed imbibition dynamics (Figure 7b) that could not be easily fitted with a Lucas-Washburn equation; this is consistent with the previous remarks, given the fact that the Vein sample only contained one row of microchannels that spanned nearly the entire length of the sample; as such, its dynamics cannot be accounted for with an averaged, effective medium approach as the detailed dynamics of filling of a unit cell accounts for most of the observed dynamics. Notably, the Vein sample started with much slower dynamics than all other samples, even Blank. We hypothesize that this slow initial propagation occurred because the very long microchannel acted as a sink for water molecules, and the observed delay corresponded to the time required to saturate the porous area below the microchannel. In the Supplemental Material, we show that we expect the timescale of such a process to be $\tau\sim 2dL_{\mathrm{cell}}/\omega_{0}$, with $L_{\mathrm{cell}}=\ell+2d$ the length of the unit cell, and $\omega_{0}$ the Lucas-Washburn coefficient of the nanoporous layer (see Theory, Equation 22). For the porous silicon that we have used in our samples, and with an applied humidity of $p/p_{\mathrm{sat}}=0.93$, we expect $\omega_{0}\simeq 5\times 10^{-9}\,\mathrm{m^{2}/s}$ [21]; with the dimensions of the Vein sample (see Table 1); with these values, we estimate $\tau\sim 25$ min, consistent with the dynamics observed in Figure 7b. In comparison, similar estimates with the other samples with microstructures (Square, Short, Long) yield $\tau<20$ s, much shorter than the observed imbibition dynamics in Figure 7b. As a result, we do not expect the filling time of individual rows of cells to be limiting for these other geometries.
@@ Appendix A Sample pictures
Figure 9 presents micrographs of the different samples used in our study.
@@ Appendix B High pressure experiment
An example of an experiment with high pressure-driven flow is shown in Figure 10.

### v3

# Present address: ]Universite Claude Bernard Lyon 1, CNRS, Institut Lumière Matière, UMR5306, F69100 Villeurbanne, France

@@ I Introduction
As we have presented in previous studies, synthetic systems inspired by the structure and hypothesized function of xylem provide a route to generating extreme scenarios in which the pore liquid experiences large tensions and to studying imbibition and permeation dynamics driven by these stresses [13, 14, 15]. Such synthetic xylem structures provide a basis for testing biophysical hypotheses [13, 16], interrogating nanoscale transport and thermodynamic phenomena [17], and designing approaches to heat transfer [18, 19] and separations [20].
@@ II.1 Samples
Table 1: Sample properties: dimensions of microchannel segments ($w$, $\ell$, $d$) as defined in Figure 2, and corresponding aspect ratio ($\xi$) and area coverage ratio ($\gamma$); $\kappa/\kappa_{0}$ is the permeability increase predicted by our numerical simulations (see section III, Theory), or by our approximate analytical formula (Equation 17), except for the Blank sample for which $\kappa/\kappa_{0}=1$ by definition. The Blank sample contained no microchannels such that the geometrical parameters are not applicable (n.a.) in this case.
@@
We fabricated five different geometries for the microstructure: a reference sample (Blank) without any etched microchannels in the glass, serving as a reference for the measure of transport in the bare porous silicon layer (permeability $\kappa_{0}\,\mathrm{[m^{2}/(Pa.s)]}$), and four structures of similar surface coverage ($\gamma=S_{\mathrm{microchannels}}/S_{\mathrm{poSi}}\mathrm{[-]}$, where $S_{\mathrm{microchannels}}\,\mathrm{[m^{2}]}$ is the in-plane area of all microchannels and $S_{\mathrm{poSi}}$ is the total in-plane area of the porous silicon through which transport occurred), but strongly varying in-plane aspect ratio ($\xi\mathrm{[-]}$). Table 1 provides the geometries of the features of the five samples; see also Appendix A, Figure 9 for actual micrographs of every sample. For all samples, the structure was repeated across a zone of length $L=1$ cm and width $W=2.2$ cm, except for the vein sample with $L=1.5$ cm. As a result, each vein in this latter sample spanned almost the whole length of the medium, while for other samples the pattern was repeated in the $y$ direction.
@@ II.2 Experiments
For high pressure flow experiments (Figure 3a), we first removed air from the pores and microchannels by evacuation in vacuum. Then, we placed the samples in a commercial, high-pressure vessel (HIP Inc.) 111 Note that appropriate equipment and protection must be used when handling fluids at high pressures. filled with water, with an applied pressure, $P$ (typically 3 to 20 MPa depending on the sample). This high pressure resulted in an inwards flow, progressively filling the serpentine channel and shrinking the bubble at the backend of the sample (bottom of structures in Figure 2). Samples were evacuated prior to filling to avoid the pressure associated with compressing and dissolving trapped air. We left the samples under pressure for intervals of time, $\Delta t$ before releasing the pressure. We took snapshots of the samples out of the vessel with a stereoscope and camera (Leica MZFLIII stereoscope from Leica Microsystems GmbH, Wetzlar, Germany, and a QImaging Retiga 1300 camera from QImaging, Surrey, Canada), in order to measure the displacement, $\Delta X$ of the meniscus in the serpentine channel (see Appendix B, Figure 10); we repeated this process until filling was complete. We estimated the velocity for each period of pressurization, $U=\Delta X/\Delta t$. This process was required because the pressure vessel was opaque. We then estimated the mass flow rate, Q [kg/s] as:
@@
We note that in some experiments, the meniscus was in the distribution channel rather than in the serpentine channel (see e.g. Appendix B, Figure 10). In both cases, we used the known cross-sectional area of the channel, $s\,\mathrm{[m^{2}]}$, in which we tracked the meniscus to convert the speed of the meniscus into a volumetric flow rate.
@@
In this scenario, inspired by xylem sap flow in plants, flow through the structure was driven by evaporation at the open edge (the "leaf") into a subsaturated vapor (Figure 3b). In a previous publication [17], we showed that this method provided an accurate way of measuring permeability of nanoporous samples. Here, we apply it to the composite, synthetic xylem structures. In order to perform transpiration experiments, we used samples that were completely filled with water except for a remaining vapor bubble in the serpentine channel. This situation occurred naturally at the end of a high pressure experiment (see above). We then placed the samples in a humidity-controlled environment with sub-saturated water vapor (water vapor pressure, $p<p_{\mathrm{sat}}(T)$). Due to the local equilibrium at the open edge between liquid water in the nanopores and the external water vapor (see section III, Theory), a reduced (typically negative) pressure developed at the open edge, resulting, after a short transient, in a steady-state flow through the structure [17]. This flow progressively emptied the serpentine channel; we tracked the resulting motion of the liquid/vapor meniscus in the serpentine channel, using a camera (Point Grey Grasshopper Monochrome camera) a macro lens (AF Micro-Nikkor 60 mm f/2.8), and white, diffuse illumination (Schott ACE 150W light source). From the velocity of the meniscus, we calculated the mass flow rate through the structure using Equation (1). For every sample we recorded transpiration-induced flow at a variety of imposed vapor pressures, $p$, in a range high enough to avoid triggering either dewetting (desorption) from the nanopores or cavitation within the microchannels (see section IV, Results and Discussion).
@@
We used spontaneous imbibition as a third method to drive transport within our structures. As we have shown recently [21], liquid imbibition in a nanoporous medium can occur spontaneously by condensation of the vapor into the nanopores, provided that the vapor pressure of the external vapor is above the threshold for capillary condensation, $p_{\mathrm{cond}}$. With our poSi layer, $p_{\mathrm{cond}}\simeq 0.6\,p_{\mathrm{sat}}$ [21], so that we worked with an imposed vapor pressure above this value, $p=0.93\,p_{\mathrm{sat}}$ to induce condensation and imbibition. We followed the imbibition dynamics optically, using time-lapse videos obtained with the same imaging system as for transpiration experiments (camera and macro lens, see above). Because the optical reflectance of the poSi layers changes as a function of water content (see section IV, Results and Discussion), we tracked the imbibition front by monitoring relative changes in the gray level of the image, $\Delta I/I$ averaged in the $x$ direction; we then extracted the position of the imbibition front by defining $Y(t)$ as the position at which the signal is at its midpoint between the dry gray level value and the wet value (see [21] for details). From the displacement, $Y(t)$ of the invading liquid front within the structure, we measured the dynamics of imbibition in our structured samples and estimated the effective permeability of the sample using a modified Lucas-Washburn equation (see section III, Theory).
We performed two types of imbibition experiments. For the first type, we first dried the samples for at least 48 hours in dry air, with a relative humidity (RH) of $6$%. Then, we placed the samples in a closed box with a humidity of $93$% RH, set by an unsaturated solution of sodium chloride (NaCl). For the second type, the samples were evacuated for at least 24 hours in a vacuum chamber. Then, we released water vapor in the chamber to impose a fixed vapor pressure of $p=0.93\,p_{\mathrm{sat}}$ ($93$% RH). As a result, imbibition occurred with air-filled nanopores and microchannels in the first type of experiments and with evacuated nanopores and microchannels in the second type.
@@ III.3 Relative magnitudes of transport mechanisms
$\frac{q_{\mathrm{v,vac}}}{q_{\mathrm{po}}}=\frac{(\epsilon aH_{\mathrm{v}})^{2}}{12\eta^{\ast}\kappa_{0}}=\mathcal{O}(10^{2})$ (10)
@@
$\frac{q_{\mathrm{v,air}}}{q_{\mathrm{po}}}=\frac{\epsilon^{2}aD}{p_{\mathrm{sat}}\kappa_{0}}=\mathcal{O}(10^{-1})$ (11)
@@ III.4 Effective permeability of the composite structures
Figure 5: Numerical and analytical predictions of steady-state flow through the composite matrix, as a function of the microchannel aspect ratio ($\xi=\ell/w$) and of the surface density of channels ($\gamma$). (a) Pressure field (colormap) and streamlines (black lines) for the unit cell of the composite porous medium, obtained from numerical simulations (here with $\gamma=0.3$, $\xi=2.2$); inside the white dotted line is the microchannel over the porous layer and outside this line is just the porous substrate. In this prediction, we assumed that the microchannels had infinite permeability, a reasonable approximation when the microchannels are filled with liquid water or pure water vapor (see Theory, Equations (9) and (10)). (b) Comparison of numerical and analytical predictions of the effective permeability ratio $\kappa/\kappa_{0}$ as a function of $\xi$ for different values of $\gamma$. Symbols represent predictions of the numerical simulations as in (a), and lines represent approximate, analytical predictions based on a resistors-in-series approach (Equation 18). Sketches show the appearance of a unit cell for the three different values of $\gamma$, at an aspect ratio, $\xi=8$; the data corresponding to these three sketches is circled in gray (dashes). See Appendix C, Figure 11 for a similar comparison between simulation results and a Maxwell-Garnett-like approach (Equation 16).
@@
(details of the derivation can be found in Appendix C, and in [30]). Equation (16) gave satisfactory agreement with our simulation results, but only for low areal fractions, $\gamma$, and moderate aspect ratios, $\xi$ (see Appendix C, Figure 11). This observation is not surprising, given that Zimmerman’s calculation is a first-order approach which performs best at low densities of inclusions [32]. As a result, we also used another approach better suited for large areal coverages and large aspect ratios.
@@ IV.1 High pressure flow
We used high pressure, steady-state flow in our various composite structures to estimate the permeability associated with samples containing microstructure (samples Square, Short, Long, Vein), relative to that of a purely nanoporous sample (Blank, permeability $\kappa_{0}$). From the observed filling velocity $U=dX/dt$ of the meniscus in the serpentine channel (see Methods and Appendix B, Figure 10), and Equations (1) and (19) we estimated the permeability ratio $\kappa/\kappa_{0}$, where the index 00 refers to the reference sample (Blank). We report and discuss the extracted values of the permeability ratios in the general discussion (see below, section IV.4 and Figure 8b).
@@ IV.3 Imbibition
Figure 7: Dynamics of spontaneous imbibition by capillary condensation (see Figure 3c). (a) Images recorded during the experiments in the Long sample when evacuated. The darker zone on top of the images is the area wetted by the fluid (contrast has been enhanced for clarity). See Supplemental Material for the corresponding complete video (Appendix E). (b-c) Dynamics of imbibition for all samples when evacuated (b) and air-filled (c), represented by the squared position, $Y^{2}$, of the wetting front as a function of time (see Figure 3c). Different colors represent the different sample geometries, and black lines represent linear fits, indicating Lucas-Washburn-like dynamics on average (Equation 22).
@@
Noticeably, the Vein sample displayed imbibition dynamics (Figure 7b) that could not be easily fitted with a Lucas-Washburn equation; this is consistent with the previous remarks, given the fact that the Vein sample only contained one row of microchannels that spanned nearly the entire length of the sample; as such, its dynamics cannot be accounted for with an averaged, effective medium approach as the detailed dynamics of filling of a unit cell accounts for most of the observed dynamics. Notably, the Vein sample started with much slower dynamics than all other samples, even Blank. We hypothesize that this slow initial propagation occurred because the very long microchannel acted as a sink for water molecules, and the observed delay corresponded to the time required to saturate the porous area below the microchannel. In Appendix D, we show that we expect the timescale of such a process to be $\tau\sim 2dL_{\mathrm{cell}}/\omega_{0}$, with $L_{\mathrm{cell}}=\ell+2d$ the length of the unit cell, and $\omega_{0}$ the Lucas-Washburn coefficient of the nanoporous layer (see Theory, Equation 22). For the porous silicon that we have used in our samples, and with an applied humidity of $p/p_{\mathrm{sat}}=0.93$, we expect $\omega_{0}\simeq 5\times 10^{-9}\,\mathrm{m^{2}/s}$ [21]; with the dimensions of the Vein sample (see Table 1); with these values, we estimate $\tau\sim 25$ min, consistent with the dynamics observed in Figure 7b. In comparison, similar estimates with the other samples with microstructures (Square, Short, Long) yield $\tau<20$ s, much shorter than the observed imbibition dynamics in Figure 7b. As a result, we do not expect the filling time of individual rows of cells to be limiting for these other geometries.
@@ Appendix A Sample pictures
For the present study, we have used five different samples, the properties of which being listed in Table 1. In Fig. 9, we also provide actual micrographs of all samples. The micrographs for the square and long samples are identical to those in Fig. 2b.
@@ Appendix B High pressure experiment
In sections II.2 and IV.1 we have described high pressure flow experiments, from which the permeability of the composite structures can be estimated. In Fig. 10, we have provided an example of a partial image sequence obtained in one of these experiments.